log[5](x-2)+log[5](2x-9)=2
Answer by RAY100(1637)

log(5){x-2} +log(5){2x-9} = 2
.
log a + log b = log a*b
.
log(5) { (x-2)(2x-9)} = 2
.
log(a) b = log b / log a
.
log{ (x-2)(2x-9)} / log(5) = 2
.
log{ 2x^2 -9x -4x +18} = 2 log 5 = 1.3979
.
Raise each side as power of 10,,,,10^(log a) =a
.
{ 2x^2 -13x +18} = 10^1.3979 = 25
.
2x^2 -13x -7 =0,,,,,,,,factor
.
(2x+1)(x-7) =0
.
x = 7,,,, - 1/2 ( not realistic)
.
check,,(x=7),,,,,,log(5){5} +log(5){5} = 2
.
log 25 / log 5 = 2
.
log 25 =2 log 5 = 1.3979,,,,,,ok
.